OpenStudy (lalaland_lauren):

Determine the factors of 15x2 + 3xy + 10x + 2y. (3x + 2)(5x + y) (5x + y)(2x + 3) (3x + y)(5x + 2) (5x + 3)(2x + y)

OpenStudy (lalaland_lauren):

@satellite73

OpenStudy (lalaland_lauren):

@vmac2122

OpenStudy (misty1212):

HI!!

OpenStudy (lalaland_lauren):

Hello :)

OpenStudy (misty1212):

we can multiply each one out and see what works, or perhaps we can work smarter

OpenStudy (misty1212):

you see how the last term is \(2y\) right?

OpenStudy (lalaland_lauren):

That's my problem, I get so confused multiplying everything out :( and yeah

OpenStudy (misty1212):

that gets rid if b and d do you see why?

OpenStudy (lalaland_lauren):

Yeah, that's actually a really cool loophole

OpenStudy (misty1212):

\[ (5x + y)(2x + 3)\] last term will be \(3y\) so that is no good

OpenStudy (misty1212):

same here \[ (5x + 3)(2x + y)\]

OpenStudy (lalaland_lauren):

I figured it out, it's A

OpenStudy (misty1212):

so we can try the first one \[ (3x + 2)(5x + y)\]

OpenStudy (misty1212):

yeah that looks good

OpenStudy (lalaland_lauren):

Yayyyyy I got it! :)

OpenStudy (misty1212):

\[\huge \color\magenta\heartsuit\]

OpenStudy (lalaland_lauren):

Thank youuuu!

OpenStudy (misty1212):

you are quite welcome got more?

OpenStudy (lalaland_lauren):

A few, yes hold on

OpenStudy (misty1212):

kk

OpenStudy (lalaland_lauren):

Graph of a cubic polynomial that falls to the left and rises to the right with graph touching x axis at negative 1 and x intercept at 1 Which of the following functions best represents the graph? f(x) = x3 + x2 − 4x − 4 f(x) = x3 + x2 − x − 1 f(x) = x3 + 3x2 − 4x − 12 f(x) = x3 + 2x2 − 2x − 1

OpenStudy (misty1212):

ok we can do this

OpenStudy (misty1212):

you see how it only touches where \(x=-1\)?

OpenStudy (lalaland_lauren):

Yeah, and where x=1

OpenStudy (misty1212):

actually there is a difference, where \(x=1\) it CROSSES the axis, at \(x=-1\) it comes up, touches it, but does not cross, goes back down

OpenStudy (misty1212):

so it has to zeros, one at \(x=-1\) and one at \(x=1\) because it crosses at \(x=1\) one factor is \((x-1)\)

OpenStudy (misty1212):

because it touches at \(x=-1\) then the factor looks like \((x+1)^2\)

OpenStudy (misty1212):

lol "has two zeros" i meant

OpenStudy (misty1212):

final job is to multiply \[(x+1)^2(x-1)\] to see what we get i would cheat!!

OpenStudy (misty1212):

looks like \[x^3+x^2-x-1\] is a winner is that a choice?

OpenStudy (lalaland_lauren):

Yup :) Wolfram Alpha has been helping me a lot this unit lol

OpenStudy (misty1212):

might as well use it instead of making an algebra mistake i do!

OpenStudy (lalaland_lauren):

Haha! I have just one more question and them I'm done lol

OpenStudy (lalaland_lauren):

Which shows 602 − 402 being evaluated using the difference of squares method? 602 − 402 = (3600 + 1600)(3600 − 1600) = 10,400,000 602 − 402 = 3600 − 1600 = 2,000 602 − 402 = (60 − 40)2 = 202 = 400 602 − 402 = (60 + 40)(60 − 40) = (100)(20) = 2,000

OpenStudy (lalaland_lauren):

The twos are squares

OpenStudy (misty1212):

\[60^2-40^2\]?

OpenStudy (lalaland_lauren):

Wait nvm I found the answer haha

OpenStudy (misty1212):

D right?

OpenStudy (lalaland_lauren):

Yup xD

OpenStudy (lalaland_lauren):

I got an 88% :D

OpenStudy (misty1212):

that it?

OpenStudy (misty1212):

then time for a tune https://www.youtube.com/watch?v=smdNetoPFV0

OpenStudy (lalaland_lauren):

I LOVE THAT SONG!!!!!

OpenStudy (lalaland_lauren):

I love that whole band, actually!

OpenStudy (misty1212):

lol thought so !!

OpenStudy (lalaland_lauren):

How'd you know? XD

OpenStudy (misty1212):

american psychic

OpenStudy (lalaland_lauren):

*clapping emoji*

OpenStudy (misty1212):

\[\color\red\heartsuit\]

OpenStudy (lalaland_lauren):

Unfortunately I have to go to bed now :( I'm extremely tired and I'm pretty sure I have to get up early to keep working on assignments. Will you be on in the morning?

OpenStudy (misty1212):

i can be here at about 9 or 9:30 earlier probably not

OpenStudy (misty1212):

tag me if you like, i will try to help if i can!

OpenStudy (misty1212):

my bed time too

OpenStudy (lalaland_lauren):

That's a good time for me :) alrighty, thanks so much for helping me! good night c:

OpenStudy (misty1212):

gnight c: